OpenStudy (anonymous):

write the equation of the line in slope intercept form that is parallel to 6x-3y=-9 and contains the points (-5 -2) please help !

OpenStudy (anonymous):

find slope of 6x-3y=9 then use y-(-5)=slope* (x-(-2))

OpenStudy (anonymous):

is it 23 ? @psi9epsilon

OpenStudy (anonymous):

slope is 2 6x-3y=9 is same as 3y=6x-9 y=2x-3 compare this y=mx+c where m=slope

OpenStudy (anonymous):

how would i make it parallel though ? @psi9epsilon

OpenStudy (anonymous):

it says that the line is parallel, hence they have same slope hope that helps

OpenStudy (anonymous):

oh okay, thanks though